If 6m = 156, then m = what.
Answer by TimothyLamb(4379)

---
6m = 156
---
divide both sides by 6
---
6m/6 = 156/6
(6/6)m = 156/6
(6/6) = 1
---
m = 156/6
m = 26
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php